# Consistent estimation of generative model representations in the data kernel perspective space

### v1

@@ 5.2 Fixed collection of models and growing set of queries
We next consider the setting where $n=10$ is fixed and both $m$ and $r$ grow. We use the same $10$ models as above. The target DKPS $\widehat{\boldsymbol{\psi}}^{*}$ is shown on the left of Figure 2 and was estimated with $m=200$ and $r=50$. As with Figure 1, the red circles have radius equal to the average (across bootstrap samples) $L_{2}$ norm of the difference between $\widehat{\boldsymbol{\psi}}^{*}$ and $\widehat{\boldsymbol{\psi}}\mathbf{W}$ for selected $(m,r)$ pairs. The right figure of Figure 2 shows the average two-to-infinity norm of the difference between $\widehat{\boldsymbol{\psi}}^{*}$ and $\widehat{\boldsymbol{\psi}}$ across bootstrap samples for more values of $(m,r)$. The two-to-infinity norm decreases as both $(m,r)$ increase, which supports Theorem 3.
The right figure of Figure 2 shows the possibility of a computational trade-off between increasing $m$ and increasing $r$. In particular, the computational cost of each iteration of the experiment is approximately $O(mr)$: getting more replicates of a fixed number of queries is approximately the same as getting a small set of replicates for more queries from a compute stand point. From an estimation stand point, however, the right figure of Figure 2 shows that getting more queries may be more beneficial. Per Theorem 2, this observation depends on the distributional properties of each $F_{ij}$.
Figure 2: Numerical evidence of the consistency of $\widehat{\boldsymbol{\psi}}$ to $\boldsymbol{\psi}$ for fixed $n=20$ and growing $m$. The left figure shows the 2-d DKPS of $20$ models induced with $200$ queries and $50$ replicates each. The red circles have radius equal to the average $L_{2}$ norm between the “ground truth" and model representations estimated for selected $(m,r)$ pairs. The right figure shows the average maximum row $L_{2}$ norm for various $(m,r)$ pairs.
@@ 5.3 Growing collection of models and growing set of queries
Finally, we consider the setting with growing $n$ and growing $m$. The target DKPS $\widehat{\boldsymbol{\psi}}^{*}$ was estimated using all $50$ models, all $200$ queries and all $50$ replicates per query. For each $(n,m,r)$ triple, we randomly sample $n$ models, $m$ queries, and $r$ replicates without replacement and estimate the DKPS $\widehat{\boldsymbol{\psi}}$. Figure 3 shows the average two-to-infinity norm of the difference between $\widehat{\boldsymbol{\psi}}^{*}$ and $\widehat{\boldsymbol{\psi}}\mathbf{W}$ across bootstrap samples for the $n$ sampled models. As each element of the triple increases, the reported difference gets close to 00, which supports Theorem 5.
@@ 6 Discussion
In [12], a novel method was proposed for embedding generative models into a finite-dimensional Euclidean space in the context of queries – the perspective space. In this paper, we analyzed the estimated perspective space in regimes where the number of models and/or the number of queries can remain fixed or grow and demonstrated that it is consistent for a population-level perspective space. Importantly, we also describe sufficient conditions for consistent estimation of the perspective space. In this regard, we establish that if the number of queries and/or models grow adequately slowly with the number of replicates relative to distributional properties of the $F_{ij}$ then the perspective space can be estimated consistently.
Low-dimensional representations of collections of generative models enable the use of classical methods to understand differences in model behavior and, eventually, to make sense of model evolution. For example, [12] uses the perspective space to demonstrate that the communication structure underlying a system of interacting language models has an impact on system-level and model-level evolution. Going further, we could use the perspective space for entire classes of model-level inference problems such as predicting the the pre-training data mixture, predicting model safety, predicting the model’s score on a benchmark, etc. Deeper investigations into these applications are warranted and our results support the perspective space as a principled foundation for these pursuits.
Figure 3: Numerical evidence of the consistency of $\widehat{\boldsymbol{\psi}}$ to $\boldsymbol{\psi}$ for growing $n$ and growing $m$. The target DKPS $\widehat{\psi}$ was estimated using $n=50$, $m=200$, and $r=50$.
@@
Further, the consistency of the estimated perspectives is the least we could ask for [2]. We expect concentration inequalities and distributional results for the estimated perspectives to be important theoretical extensions to support applications in non-asymptotic regimes. For example, uniform convergence of the perspectives enable use of the estimated perspective space as a principled substitute for its population counterpart after a particular amount of models, queries, and replicates. Similarly, establishing distributional properties of the estimated perspectives, such as asymptotic normality, in particular settings as to practicable inference methods.
@@
We note that in most generative models in practice such as large language models with a maximum context size and maximum output sequence length and diffusion models that output RBG images, there are only a finite many possible input and output sequences or images. Hence, the maximum trace of the covariance matrices of $\{F_{ij}\}$ does not grow without bound and, importantly, we are in the regime where $m$ or $n$ can grow arbitrarily fast with respect to $r$.
Lastly, our results are for representations of models in the context of a set of queries. There are other potential vector representations of models that do not depend on a set of queries, such as model weights. The geometry of these quantities can similarly be obtained in settings where an appropriate dissimilarity is defined. In our setting, different sets of queries will approximate the non-query dependent geometry to different degrees. Investigations into how well different query sets approximate more inherent features of the models are necessary, as well as investigations into trade-offs between query-independent and query-dependent model representations.
@@ A.2 Fixed set of models and growing set of queries
$\lim_{r\to\infty}\frac{\frac{1}{m}\sum_{j=1}^{m}\gamma_{ij}}{r}=0$
Proof. Obvious consequence of Theorem 2 and Theorem 3.
@@ A.3 Growing set of models and growing set of queries
Proof. Same argument as Theorem 3.
@@
Proof. Obvious consequence of Lemma 2 and Theorem 4.

### v2

@@ 5.2 Fixed collection of models and growing set of queries
We next consider the setting where $n$ is fixed and both $m$ and $r$ grow. We use the same models as above. The two target DKPS $\widehat{\boldsymbol{\psi}}^{*}$ are shown on the left for each pair of Figure 2. They were estimated with $m=M$ and $r=R$. As with Figure 1, the red circles have radius equal to the average (across bootstrap samples) $L_{2}$ norm of the difference between $\widehat{\boldsymbol{\psi}}^{*}$ and $\widehat{\boldsymbol{\psi}}\mathbf{W}$ for selected $(m,r)$ pairs. The right figure of Figure 2 shows the average two-to-infinity norm of the difference between $\widehat{\boldsymbol{\psi}}^{*}$ and $\widehat{\boldsymbol{\psi}}$ across bootstrap samples for more values of $(m,r)$. The two-to-infinity norm decreases as both $(m,r)$ increase, which supports Theorem 3.
The right figure of each pair figures of Figure 2 shows the possibility of a computational trade-off between increasing $m$ and increasing $r$. In particular, the computational cost of each iteration of the experiment is approximately $O(mr)$: getting more replicates of a fixed number of queries is approximately the same as getting a small set of replicates for more queries from a compute stand point. From an estimation stand point, however, the right figures of Figure 2 shows that getting more queries may be more beneficial. Per Theorem 2, this observation depends on the distributional properties of each $F_{ij}$.
@@ 5.3 Growing collection of models and growing set of queries
Figure 3: Numerical evidence of the consistency of $\widehat{\boldsymbol{\psi}}$ to $\boldsymbol{\psi}$ for growing $n$ and growing $m$. The two target DKPS $\widehat{\psi}$ were estimated using $N$, $M$, and $R$.
Finally, we consider the setting with growing $n$ and growing $m$. The two target DKPS $\widehat{\boldsymbol{\psi}}^{*}$ were estimated using all $N$ models, all $M$ queries and all $R$ replicates per query. For each $(n,m,r)$ triple, we randomly sample $n$ models, $m$ queries, and $r$ replicates with replacement and estimate the DKPS $\widehat{\boldsymbol{\psi}}$. Figure 3 shows the average two-to-infinity norm of the difference between $\widehat{\boldsymbol{\psi}}^{*}$ and $\widehat{\boldsymbol{\psi}}\mathbf{W}$ across bootstrap samples for the $n$ sampled models. As each element of the triple increases, the reported difference gets close to 00, which supports Theorem 5.
@@ 6 Discussion
In [13], a novel method was proposed for embedding generative models into a finite-dimensional Euclidean space in the context of queries – the data kernel perspective space. In this paper, we analyzed the estimated perspective space in regimes where the number of models and/or the number of queries can remain fixed or grow and demonstrated that it is consistent for a population-level perspective space. Importantly, we also describe sufficient conditions for consistent estimation of the perspective space. In this regard, we establish that if the number of queries and/or models grow adequately slowly with the number of replicates relative to distributional properties of the $F_{ij}$ then the perspective space can be estimated consistently.
Low-dimensional representations of collections of generative models enable the use of classical methods to understand differences in model behavior and, eventually, to make sense of model evolution. For example, [13] uses the perspective space to demonstrate that the communication structure underlying a system of interacting language models has an impact on system-level and model-level dynamics. Going further, the perspective space can be used for model-level inference problems such as predicting the the pre-training data mixture, predicting model safety, predicting the model’s score on a benchmark, etc. Deeper investigations into these applications are warranted and our results support the perspective space as a principled foundation for these pursuits.
@@
Further, the consistency of the estimated perspectives is the least we could ask for [2]. We expect concentration inequalities and distributional results for the estimated perspectives to be important theoretical extensions to support applications in non-asymptotic regimes. For example, establishing uniform convergence of the perspectives would enable use of the estimated perspective space as a principled substitute for its population counterpart after a particular amount of models, queries, and replicates. Similarly, establishing distributional properties of the estimated perspectives, such as asymptotic normality, in particular settings as to practicable inference methods.
@@
We note that for most generative models in practice such as large language models with a maximum context size and maximum output sequence length and diffusion models that output RBG images, there are only a finite many possible input and output sequences or images. Hence, the maximum trace of the covariance matrices of $\{F_{ij}\}$ does not grow without bound and, importantly, we are in the regime where $m$ or $n$ can grow arbitrarily fast with respect to $r$.
Lastly, our results are for representations of models in the context of a set of queries. There are other potential vector representations of models that do not depend on a set of queries, such as model weights. The geometry of these quantities can similarly be obtained in settings where an appropriate dissimilarity is defined. In our setting, different sets of queries will approximate the non-query dependent geometry to different degrees. Investigations into how well different query sets approximate more inherent features of the models are necessary, as well as investigations into practical trade-offs between query-independent and query-dependent model representations.
@@ A.2 Fixed set of models and growing set of queries
Theorem 3. In the setting of Theorem 2, suppose for all $i\in[n]$, $\frac{1}{m}\sum_{j=1}^{m}\gamma_{ij}=o(r)$. Also, let $\widehat{\boldsymbol{\psi}}\in\mathrm{MDS}(\mathbf{D})\subset\mathbb{R}^{n\times d}$. There exists a subsequence $\{r_{u}\}_{u=1}^{\infty}$ of $\{r\}_{r=1}^{\infty}$ such that for all $(i,i^{\prime})\in[n]^{2}$,
Proof. Let us assume $\frac{1}{m}\sum_{j=1}^{m}\gamma_{ij}=o(r)$. From Theorem 2, then we must have $\left\lVert\mathbf{D}-\boldsymbol{\Delta}^{(\infty)}\right\rVert_{F}\to^{P}0$ as $r\to\infty$. From Lemma 1, this means there exists a subsequence $\{r_{u}\}_{u=1}^{\infty}$ of $\{r\}_{r=1}^{\infty}$ such that for all pairs $(i,i^{\prime})\in[n]^{2}$,
$\left(\left\lVert\widehat{\boldsymbol{\psi}}_{i}^{(r_{u})}-\widehat{\boldsymbol{\psi}}_{i^{\prime}}^{(r_{u})}\right\rVert-\left\lVert\boldsymbol{\psi}_{i}-\boldsymbol{\psi}_{i^{\prime}}\right\rVert\right)\to^{P}0$
as $u\to\infty$.
@@ A.3 Growing set of models and growing set of queries
Proof. Recall (from the proof of Theorem 2) that $\left\lvert\mathbf{D}_{ii^{\prime}}-\boldsymbol{\Delta}_{ii^{\prime}}\right\rvert\leq\frac{1}{m}\left\lVert\bar{\mathbf{X}}_{i}-\boldsymbol{\mu}_{i}\right\rVert_{F}+\frac{1}{m}\left\lVert\bar{\mathbf{X}}_{i^{\prime}}-\boldsymbol{\mu}_{i^{\prime}}\right\rVert_{F}$ and that for all $i$, $\mathbb{P}\left[\frac{1}{m}\left\lVert\bar{\mathbf{X}}_{i}-\boldsymbol{\mu}_{i}\right\rVert_{F}>\epsilon\right]\leq\frac{\sum_{j=1}^{m}\gamma_{ij}}{rm\epsilon^{2}}$ for any $\epsilon>0$. Thus, if $\lim_{r\to\infty}\frac{\frac{1}{m}\sum_{j=1}^{m}\gamma_{ij}}{r}=0$ for all $i$, then $\frac{1}{m}\left\lVert\bar{\mathbf{X}}_{i}-\boldsymbol{\mu}_{i}\right\rVert\to^{P}0$ for all $i$, which in turn implies $|\mathbf{D}_{ii^{\prime}}-\boldsymbol{\Delta}_{ii^{\prime}}|\to^{P}0$ for all pairs $(i,i^{\prime})\in\mathbb{N}^{2}$ as $r\to\infty$. Using Assumption 2 and Triangle Inequality, we have that for all pairs $(i,i^{\prime})\in\mathbb{N}^{2}$, $|\mathbf{D}_{ii^{\prime}}-\boldsymbol{\Delta}^{(\infty)}(\boldsymbol{\phi}_{i},\boldsymbol{\phi}_{i^{\prime}})|\to^{P}0$ as $r\to\infty$. Using Lemma 2, for all $p\geq 1$, for some subsequence $\{r_{u}\}_{u=1}^{\infty}$ of $\{r\}_{r=1}^{\infty}$,
@@
Proof. Let us assume that $\frac{1}{m}\sum_{j=1}^{m}\gamma_{ij}=o(r)$. Then, by Theorem 4, for all pairs $(i,i^{\prime})\in\mathbb{N}^{2}$, $|\mathbf{D}_{ii^{\prime}}-\boldsymbol{\Delta}^{(\infty)}(\boldsymbol{\phi}_{i},\boldsymbol{\phi}_{i^{\prime}})|\to^{P}0$ as $r\to\infty$. Using Lemma 2, for some subsequence $\{r_{u}\}_{u=1}^{\infty}$ of $\{r\}_{r=1}^{\infty}$, for all $p\geq 1$,